OpenStudy (anonymous):

what is 7p+5=82??? xxx

OpenStudy (anonymous):

7p=82-5 7p=77 p=11

OpenStudy (anonymous):

7p + 5 = 82 7p = 82 - 5 7p = 77 p = 77/7 p = 11

OpenStudy (katrinakaif):

Can you demonstrate and guide her step by step rather than posting the solution itself?

OpenStudy (anonymous):

Thankyou everyone i will get a good grade on my gsce know :D xxxxxxx